### v1

# SimGen: Simulator-conditioned Driving Scene Generation

Abstract
Controllable synthetic data generation can substantially lower the annotation cost of training data in autonomous driving research and development. Prior works use diffusion models to generate driving images conditioned on the 3D object layout. However, those models are trained on small-scale datasets like nuScenes, which lack appearance and layout diversity. Moreover, the trained models can only generate images based on the real-world layout data from the validation set of the same dataset, where overfitting happens. In this work, we introduce a simulator-conditioned scene generation framework called SimGen that can learn to generate diverse driving scenes by mixing data from the simulator and the real world. It uses a novel cascade diffusion pipeline to address challenging sim-to-real gaps and multi-condition conflicts. A driving video dataset DIVA is collected to enhance the generative diversity of SimGen, which contains over 147.5 hours of real-world driving videos from 73 locations worldwide and simulated driving data from the MetaDrive simulator. SimGen achieves superior generation quality and diversity while preserving controllability based on the text prompt and the layout pulled from a simulator. We further demonstrate the improvements brought by SimGen for synthetic data augmentation on the BEV detection and segmentation task and showcase its capability in safety-critical data generation. Code, data, and models will be made available at https://metadriverse.github.io/simgen/.
Figure 1: We propose SimGen, a controllable scene generation paradigm through the simulator-conditioned generative model. It learns from real-world and simulated data and then can generate diverse driving scenes based on the simulator’s control conditions and text prompt.
This work is done during YZ’s visit to the University of California, Los Angeles.Corresponding author.
1 Introduction
A high-quality and diverse training data corpus is crucial for autonomous driving research and development. However, it is costly and laborious to annotate the data. Synthetic data generation is a promising alternative to harvest annotated training data, which brings realistic images and notable performance improvements across tasks like object detection [9] and semantic segmentation [73]. Besides the realism of the generated images, there are two necessary conditions for a practical synthetic data generator for autonomous driving: 1) Appearance diversity, which ensures the synthetic data can cover a spectrum of weather, environmental, and geographical conditions. 2) Layout diversity, namely the distribution of objects should cover different traffic scenarios, including safety-critical situations that are rare to collect in the real world.
Recent diffusion-based generative models show promising results to generate realistic driving images from text prompts [76], BEV road maps [65], and object boxes [18, 68, 70, 77]. Despite generating coherent images, these attempts lack the generalizability of generating new and diverse real-world appearances and traffic scenarios due to data limitations. They are confined to learning on small-scale datasets [27, 30, 42, 69] with limited scenarios such as only urban streets [5] or restricted weather conditions [50]. In addition, the driving behaviors in the available driving datasets like nuScenes are tedious and lack complex or safety-critical situations. Another option for collecting synthetic data is from driving simulators, which can effortlessly generate scenes encompassing various behaviors with its physics and graphics engines [14, 34, 56, 59, 64]. Simulators also provide accurate control over all objects and their spatial locations, thus can easily generate a huge amount of traffic layout maps. However, open-source simulators usually only contain a limited amount of 3D assets, and they lack a realistic visual appearance. Thus, the models trained on simulator-generated data can easily overfit, also known as the Simulation to Reality (Sim2Real) gap.
We take the best of two worlds by integrating the data-driven generative models with a simulator to obtain both the appearance diversity of real-world data and the layout controllability of simulated data. To this end, we introduce SimGen, a simulator-conditioned diffusion model, which follows the layout guidance from the simulator and rich text prompts to generate diverse driving scene images. One naïve approach is to guide an image generation model with the depth and semantic images from the simulator via training a control branch through ControlNet [79]. Yet, as the simulator has limited assets and cannot fully capture the variations in the real world, the simulated conditions and the underlying real-world conditions that guide a diffusion model to generate real-world images might have conflicts. To tackle this, SimGen adopts a cascade design. The model first injects noise-added simulated conditions such as depth and semantic images into the intermediate sampling process of a pre-trained text-to-real-condition diffusion network. The network then converts simulated conditions into more realistic conditions via continuous denoising, free of additional training on simulated conditions beyond this diffusion network. After that, a second diffusion module utilizes an adapter to integrate multimodal conditions and uses masks to filter conflicting data. SimGen thus achieves outstanding generation quality and diversity while preserving layout controllability by connecting with the simulator.
We construct a new dataset called DIVA to increase the appearance and layout diversity of the training data. DIVA comprises two parts: the web data and the synthesized data from the simulator. On the one hand, web data covers a worldwide range of geography, weather, scenes, and traffic elements, preserving the appearance diversity of a wide range of traffic participants. We design a data curation pipeline to collect and label YouTube driving videos. On the other hand, virtual driving videos with the traffic flow replayed from trajectory datasets or generated by a safety-critical scenario generator [78] are collected from a driving simulator [34]. In short, DIVA dataset blends real-world appearances and virtual layouts, consisting of 147.5 hours of Diverse In-the-wild and Virtual driving datA.
We summarize our contributions as follows: 1) a novel controllable image generation model SimGen incorporating a driving simulator to generate realistic driving scenarios with appearance and layout diversity; 2) a new dataset DIVA containing massive web and simulated driving videos that ensures diverse scene generation and advances simulation-to-reality research; 3) SimGen improves over counterparts like BEVGen [65], MagicDrive [18], Panacea [70], DrivingDiffusion [36], i.e., in terms of image quality, diversity, and controllability of scene generation.
2 Appearance Diversity and Layout Diversity from DIVA Dataset
We introduce a large-scale DIVA dataset containing diverse driving scenes in the real world and the simulation. It facilitates the training of generative models and tackles the simulation-to-reality (Sim2Real) challenge. Sec. 2.1 displays the statistics, composition, and annotation of the data, which comprises about 147.5 hours of driving videos. The data is collected from a vast corpus of high-quality YouTube driving videos and simulation environments in the MetaDrive simulator [34]. We use DIVA-Real and DIVA-Sim to denote the web data downloaded from YouTube and the data from the MetaDrive simulator, respectively. Comparisons with other datasets, license, and privacy considerations are detailed in Appendix B.
2.1 DIVA-Real: Appearance Diversity in Web Data
Collecting web videos. As shown in Fig. 2 (left), to streamline the process and minimize manual effort, we begin by searching for relevant keywords on YouTube to identify a batch of driving video channels. The videos are downloaded from these identified YouTube channels. We filter out unsuitable videos based on their length and resolution and proceed to download the appropriate ones. This yields hundreds of first-person driving videos, each with an average duration of one hour. Next, we sample the videos into frames at 10Hz, excluding the initial and final 30 seconds to eliminate user channel information. This process yields over 4.3 million frames, awaiting further data cleaning.
Table 1: Comparing DIVA with relevant datasets on scale, diversity, and annotations. : perception subset. +: including procedural generation [34] and safety-critical [78] data. Cts: countries; Seg: segmentation; Virt: virtual image.
Dataset Time Frames Cts. Cities Annotations
(hours) Text Depth Seg. Virt.
KITTI [20] 1.4 15k 1 1
CityScapes [12] 0.5 25k 3 50
Waymo [63] 11 390k 1 3
Argoverse 2 [72] 4.2 300k 1 6
nuPlan [7] 120 4.0M 2 4
Honda-HAD [29] 32 1.2M 1 -
nuScenes [6] 5.5 241k 2 2
DIVA-Real 120 4.3M 19 71
DIVA-Sim 27.5+ 998k+ 3 5
DIVA (All) 147.5 5.3M 22 76
Data cleaning and autolabeling. Data cleaning is vital for ensuring data quality, but manual inspection of each image is impractical. Inspired by [76], we implement an automated data-cleaning workflow to expedite the process. With the remarkable image understanding capabilities of the vision-language model (VLM), i.e. LLaMA-Adapter V2 [17], we are able to conduct the quality checks via VLM with a checklist including criteria such as non-front view, video transition, black screens, etc, to identify nonconforming images. Driving videos are chunked into five-frame batches. For each batch, the VLM chooses and assesses a random image; if this single image fails to pass checks, the entire batch of five frames will be discarded. In the autolabeling process, pre-trained models for various tasks, including BLIP2-flant5 [54], ZoeDepth [3], and Segformer [74] , are used to generate annotations of text, depth, and semantic segmentation, respectively. Eventually, over 120 hours of driving videos with rich annotations are collected.
2.2 DIVA-Sim: Layout Diversity from the Simulator
Figure 2: Construction of DIVA dataset. DIVA-Real (left) comprises driving videos collected from YouTube. We apply a Vision Language Model to filter out noisy images via a checklist and utilize off-the-shelf models to annotate text, depth, and semantic labels. Meanwhile, DIVA-Sim (right) employs scene records and control policies in a simulator to create map elements and objects. It can generate digital twins of real-world data and safety-critical scenes. Then various kinds of sensors placed in the simulation produce multimodal images. Ren.:rendered; T.D.: top-down view. Numbers and letters indicate the sequence of processes.
Simulators are capable of faithfully reconstructing real-world scenes and hence obtaining training data with layout diversity. Also, after loading the driving scenarios such as map topology from the dataset, the simulator allows changing the motions and states of the traffic participants with pre-defined rules or interactive policies that differ from the original ones. This inspires us to build Sim2Real data from the simulator. The Sim2Real data is induced from the same real-world scenarios, in which we can obtain real-world map topology, layout, and raw sensor data. At the same time, we can reconstruct the paired data from those scenarios but with reconstructed sensor data and even with altered layout and traffic flows. DIVA-Sim utilizes the MetaDrive simulator [34] and ScenarioNet [35] to gather 5.5 hours of virtual driving videos from nuScenes layouts [6] and another 22 hours from procedurally generated behaviors. It includes a set of safety-critical driving data through interactions introduced by an adversarial traffic generation method [78], further improving the diversity of our dataset.
Scene layout construction. We utilize ScenarioNet [35] to transform scenes into a unified description format suitable for simulators, known as scene records, logging map elements and objects. As illustrated by the example scene in Fig. 2 (right), loading scene records, MetaDrive [34] can reconstruct roads, blocks, and intersections, and place corresponding 3D models like vehicles, cyclists, and pedestrians, based on the recorded positions and orientations. We will reasonably select representative 3D models based on the category and dimensions of the objects. And the model’s shape is scaled based on the real dimensions to replicate the objects in the nuScenes dataset accurately. By doing so, the digital twin scenario can be faithfully reconstructed in the simulator.
Obtaining images via trajectory replay and rendering pipeline. The control policy determines the motion dynamics, while the sensors generate multimodal image data at any desired location. To create nuScenes digital twins, ReplayPolicy is applied to replay logged trajectories of all objects. Our cameras are placed in the exact pose of the nuScenes front camera, with the camera’s field of view adjusted to match that of nuScenes closely. The camera attribute can be set to multiple types to obtain a variety of sensor data. In summary, we can obtain the following conditions through the simulator: rendered RGB, depth, semantic segmentation, instance segmentation, and top-down views.
Creation of safety-critical data. Besides building digital twins of the real-world data, we can harness the simulator to continue growing the safety-critical data and enhance layout diversity. We apply the CAT method [78] to generate safety-critical data based on real-world scenarios. Specifically, we first randomly sample one scenario from the Waymo Open dataset [63]. A traffic vehicle is perturbed to attempt colliding with the ego-vehicle via adversarial interaction learning [78]. Thus, we harvest many safety-critical scenarios with adversarial driving behaviors, which might be challenging to collect in the real world. This scalable creation of the safety-critical data from the simulator is also one of the strengths of our method.
3 SimGen Framework
Figure 3: Illustration of SimGen. SimGen processes text and scene record as inputs. The text is feature-encoded and utilized in the subsequent modules, whereas the scene record undergoes a simulator rendering into simulated depth and segmentation (SimCond) and extra conditions (ExtraCond). SimCond, coupled with the text features, is fed into the CondDiff module that converts SimCond into RealCond, representing real depth and segmentation. Eventually, the text features, SimCond, and ExtraCond are inputted into the ImgDiff module, where an Adapter merges multi-source conditions into a unified control condition and generates driving scene images.
SimGen aims to generate realistic driving images based on the text prompt and the spatial conditions including semantic and depth maps from real-world datasets and the driving simulator. We incorporate a driving simulator into the data generation pipeline to achieve controllable and diverse image generation. Incorporating the simulator provides access to diverse layouts and behaviors of traffic participants, thus better closing the Sim2Real gap. However, if just conditioning the diffusion model on synthesized data from a simulator, the diffusion model will result in bad image quality due to the limited assets and the artificial rendering. We propose a cascade generative model that first transforms the simulated spatial conditions to realistic conditions as those in the dataset, then uses those realistic conditions to guide the first-view image diffusion model.
Illustrated in Fig. 3, SimGen first samples a driving scenario and a text prompt from the dataset and invokes the driving simulator MetaDrive [34] to render simulated conditions (SimCond), the synthesized depth and segmentation images. Then, the SimCond and text features are fed into a lightweight diffusion model CondDiff (Sec. 3.1) that converts simulated conditions into realistic conditions (RealCond), that resembles the real-world depth and segmentation images from YouTube and nuScenes datasets. Finally, a diffusion model called ImgDiff (Sec. 3.2) generates a driving scene according to multi-modal conditions, including RealCond, textual prompts, and optional simulated spatial conditions, including RGB images, instance maps, and top-down views, etc.
3.1 Sim-to-Real Condition Transformation
While we strive to align the simulator settings with real data, such as intrinsic and extrinsic parameters of the camera, there is still a disparity between RealCond and SimCond. The disparity arises from image mismatch, inherent flaws of the 3D models, and the simulator’s lack of background details (Sec. C.1.1). Consequently, simulator conditions require transformation to closely resemble real ones. An easy solution is to use domain adaptation [46] and consider the SimCond and RealCond as different image styles. However, training a domain transfer model that can generalize to novel scenarios requires paired SimCond and RealCond data far exceeding public datasets like nuScenes. Thus, it’s necessary to have an adaptation-free approach for Sim2Real transformation without additional training on SimCond. To achieve that, we first use data from DIVA-Real to train a diffusion model, CondDiff, that generates RealCond purely from text prompts. The training does not contain data rendered from simulators. During inference, CondDiff injects noise-added SimCond into the intermediate sampling process and converts it into realistic conditions via continuous denoising.
Learning to generate conditions from text inputs. To facilitate the learning process of CondDiff, we initiate this stage with text-to-RealCond generation. Concretely, we utilize Stable Diffusion 2.1 (SD-2.1) [58], a large-scale latent diffusion model for text-to-image generation. It is implemented as a denoising UNet, denoted by $\epsilon_{\theta}$, with multiple stacked convolutional and attention blocks, which learns to synthesize images by denoising latent noise. Let $\mathbf{x}_{0}\in\mathcal{X}$ represents a latent feature from the data distribution $p(\mathbf{x})$. Starting from $\mathbf{x}_{0}$, the training process involves gradually adding noise to procedure $\mathbf{x}_{t}$ for $t\in(0,1]$ until $\mathbf{x}_{t}$ transforms into Gaussian noise, namely forward stochastic differential equation (SDE) [25]. The model is optimized by minimizing the mean-square error:
$\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}$ $\epsilon,\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathbf{x}_{0}\sim p(\mathbf{x}),$ (1)
[…]
where $\sigma_{t}$ is a scalar function that describes the magnitude of the noise $\epsilon$ at denoising step $t$, $\alpha_{t}$ is a scalar function that denotes the magnitude of the data $\mathbf{x}_{0}$, $\theta$ parameterizes the denoiser model $\epsilon_{\theta}$, $\mathbf{\epsilon}$ is the added noise, and $\mathbf{c}$ is the text condition that guides the denoising process. The learning occurs in a compressed latent space $\mathcal{X}$ instead of the pixel space [58]. During sampling, the model iteratively denoises the final step prediction from the standard Gaussian noise to generate images.
The original SD-2.1 is trained on data from various domains unrelated to the depth and semantic images in driving scenes. As depicted in the CondDiff in the upper right of Fig. 3, we fine-tune the SD-2.1 to be a text-to-RealCond model using the triplets of text, depth and segmentation data from DIVA-Real and nuScenes, with the objective of Eq. 2. After loading the SD-2.1 checkpoint, all parameters $\theta$ of the UNet are fine-tuned at this stage, while the CLIP text encoder [53] and autoencoder [16] remain frozen. The depth and segmentation data is autolabelled by a set of perception models as discussed in Sec. 2.1.
Adaptation-free sim-to-real transformation. Now, we have a model CondDiff that generates RealCond purely from text prompts. We will then use the conditions from simulator SimCond to guide the sampling process so that we can transform SimCond to RealCond. According to SDEdit [43], the reverse SDE, where the diffusion model iteratively denoises standard Gaussian noise to generate images, can start from any intermediate time. This inspires us to insert noise-added SimCond into the intermediate time of the sampling process, and the model will use them as guidance to generate RealCond with the SimCond layouts. In detail, the module first encodes the SimCond into latent space to get $\mathbf{x}^{\text{sim}}$. It selects a specific time $t_{s}\in(0,1)$ and perturbs the input $\mathbf{x}^{\text{sim}}$ using a Gaussian noise of standard deviation $\sigma^{2}_{t_{s}}$ as follows:
$\texttt{Sample}\ \mathbf{x}^{\text{noi}}\sim\mathcal{N}(\mathbf{x}^{\text{sim}};\sigma^{2}_{t_{s}}\mathbf{I}).$ (3)
The perturbing process will effectively remove low-level details like pixel information while preserving high-level cues like rough color strokes [43]. The noise-processed image $\mathbf{x}^{\text{noi}}$ seamlessly substitutes the diffusion model’s state at time $t_{s}$ during denoising. Thus, the intermediate state $\mathbf{x}_{t_{s}}=\mathbf{x}^{\text{noi}}$ serves as a guidance to solve the corresponding reverse SDE as follows:
$p_{\theta}(\mathbf{x}_{t_{s}-1}|\mathbf{x}_{t_{s}})=\mathcal{N}(\mathbf{x}_{t_{s}-1};\mu_{\theta}(\mathbf{x}_{t_{s}},t),\Sigma_{\theta}(\mathbf{x}_{t_{s}},t_{s})),$ (4)
where $\mu_{\theta}$ and $\Sigma_{\theta}$ are determined by CondDiff $\epsilon_{\theta}$. The above equation iterates until the model generates a synthesized image $\mathbf{x}_{0}$ like RealCond at $t_{s}=0$. Throughout this process, all parameters of CondDiff remain frozen, with only SimCond $\mathbf{x}^{\text{sim}}$, text $\mathbf{c}$, and noise affecting the sampling process.
3.2 Controllable Image Generation with Multimodal Conditions
Table 2: Formats of conditions. Real/SimCond: depth and segmentation; ExtraCond: rendered RGB, instance maps, and top-down views.
Dataset RealCond SimCond ExtraCond
nuScenes
DIVA-Real
DIVA-Sim
In the second stage, we will use a diffusion-based model to synthesize diverse driving images by integrating various control conditions (Sec. 3.2), including the RealCond from the data or generated from SimCond by CondDiff, the textual prompt, and some extra conditions ExtraCond such as rendered RGB, instance segmentation, and top-down views from the simulator. ExtraCond offers additional information for the output image, including road typology and object attributes (orientation, outlines, and 3D locations), highlighting the necessity of incorporating them into model control.
However, there exist conflicts among multimodal conditions (Sec. C.1.2): 1) Modal discrepancy: The nuScene dataset contains a full set of RealCond, SimCond, and ExtraCond, while YouTube only includes RealCond. This might impact the quality of images generated based on nuScenes layouts due to the data bias for diffusion models [31]. 2) Condition disparity: The lack of rich background information in simulated conditions compared to real ones results in a struggle between the two modalities. In real-world images, the background might contain urban buildings with drastically different facades and street trees of different species. Although CondDiff can convert SimCond to RealCond, the domain gap prevents the same transformation for ExtraCond (e.g., rendered RGB, instance segmentation, and top-down views) from the simulator. Thus, we propose using a unified adapter in ImgDiff to address these issues. Its essence lies in mapping variable conditions into fixed-length vectors, overcoming the misalignment of low-level features, and enabling a unified control input interface for the diffusion model.
Mitigating condition conflicts with adapters. Adapters are essential at the guiding branch of image generation to ensure the model learns necessary, unique, non-conflicting information from all conditions. Inspired by UniControl [52], we devised a set of convolutional modules as the adapters to capture features from various modalities, as shown in the ImgDiff in the lower right of Fig. 3. For a set of input conditions $\mathbf{X}=\{\mathbf{x}^{1},\mathbf{x}^{2},...,\mathbf{x}^{K}\}$, each condition undergoes feature extraction via the unified adapter $\mathcal{F}_{\text{ada}}$ represented as:
$\mathcal{F}_{\text{ada}}(\mathbf{x}^{k}):=\sum^{K}_{i=1}\mathbbm{1}_{i=k}\mathcal{F}^{(i)}_{\text{cov1}}(\mathcal{F}^{(i)}_{\text{cov2}}(\mathbf{x}^{k}\cdot\mathbf{M}^{k})),$ (5)
where $\mathbbm{1}$ is the indicator function, $\mathbf{x}^{k}$ is the $k$-th condition image, and $\mathcal{F}^{(i)}_{\text{cov1}}$, $\mathcal{F}^{(i)}_{\text{cov2}}$ are the convolution layers of the $i$-th module of the adapter. $\mathbf{M}^{k}$ is the valid mask for each condition.
The valid mask is the key to mitigating conflicts. The entire mask will be padded with 0 if a condition is missing or not provided. For simulator-generated conditions, we set the masks of backgrounds to 0 based on the semantic labels, preventing unwanted constraints on background generation. Since top-down view conditions don’t belong to the frontal perspective, all information is retained. Ultimately, two convolutional layers process the concatenated condition features, max pooling them into a fixed-length feature vector for control.
Controllable image generation. We utilize the ControlNet [79] to guide image generation. After the feature extraction by $\mathcal{F}_{\text{ada}}$, conditions are encoded into the UNet model. Then, the model injects control information into each UNet layer through residual connections. All parameters in UNet’s input and middle layers are frozen, and we only fine-tune the output layers and the control branch.
4 Experiments
Setup and Protocols. SimGen is learned in two stages on DIVA and nuScenes dataset [6]. The performance is evaluated based on image quality, controllability, and diversity. The Frame-wise Fréchet Inception Distance (FID) evaluates the synthesized data’s quality. SimGen’s controllability corresponds to how well the generated images align with ground truths from the nuScenes validation set. The controllability is measured by the 3D detection metrics (AP) and BEV segmentation metrics (mIoU) when applying out-of-the-box perception models on the generated images. Lastly, diversity is measured using the pixel variance of the generated images. More details on training, sampling, and evaluation metrics are provided in Appendix C and Appendix D.
4.1 Comparison to State-of-the-arts
Table 3: Generation quality and diversity compared to nuScenes experts. The FID and $D_{\text{pix}}$ indicate the image quality and pixel diversity, respectively. gray: main metric. bold: best results.
[…]
BEVGen [65] nuScenes 25.5 17.0
BEVControl [77] 24.9 -
MagicDrive [18] 16.6 19.7
Panacea [70] 17.0 -
DrivingDiffusion [36] 15.9 20.1
SimGen-nuSc nuScenes 15.6 20.5
SimGen DIVA 15.6 26.6
Comparison to nuScenes-specific models. We compare SimGen with the most recently available data generation approaches exclusively trained on nuScenes. Sec. 4.1 shows that SimGen surpasses all previous methods in image quality (FID) and diversity ($D_{\text{pix}}$). Specifically, SimGen significantly increases $D_{\text{pix}}$ by +6.5 compared to DrivingDiffusion [36]. For fair comparisons, we train a model variant (SimGen-nuSc) on the nuScenes dataset only. We find that although SimGen-nuSc performs on par with SimGen on nuScenes, its performance in diversity is less than ideal, and it struggles to generalize to novel appearances like Desert, Mountains, and Blizzard, where the generation degrades to the nuScenes visual pattern. In contrast, SimGen trained on DIVA exhibits strong generalization ability across appearances as shown in Fig. 4.
Figure 4: Generating diverse appearances conditioned on simulator’s conditions and texts. We show the generation results of SimGen (blue boxes) and SimGen-nuSc (gray boxes) under the same conditions. Compared to models confined to limited datasets, SimGen exhibits a stronger ability to generate more realistic and diverse driving scenarios.
Controllability for autonomous driving. The controllability of our method is quantitatively assessed based on the perception performance metrics obtained using a single-frame version of BEVFusion [38]. We feed the data from nuScenes validation set into SimGen and generate the driving images. Then, the perception performance of pre-trained BEVFusion, involving map segmentation (mIoU) and 3D object detection (AP), is recorded. Compared to the perception scores on the raw nuScenes data, the relative performance metrics serve as the indicators of the alignment between the generated images and the conditions. As depicted in Tab. 5, SimGen achieves a relative performance of -3.3 on map segmentation of vehicles, underscoring a robust alignment of the generated samples.
Data augmentation via synthetic data. SimGen can produce augmented data with accurate annotation controls, enhancing the training for perception tasks, e.g., map segmentation, and 3D object detection. For these tasks, we augment an equal number of images as in nuScenes dataset, ensuring consistent training iterations and batch sizes for fair comparisons to the baseline. Tab. 5 indicates that blending generated with real data can elevate the singe-frame version of BEVFusion’s vehicle mIoU to 39.0, a +4.4 uptick compared to models trained purely on real data. These outcomes reinforce SimGen’s validity as a controllable synthetic data generator for enhancing perception models.
Table 4: Generation controllability for perception tasks. Oracle: a single-frame version of BEVFusion [38]. In blue is the relative drop compared to standard nuScenes validation data.
Method Map Seg Object Detection
$\text{mIoU}_{\text{Road}}$ $\text{mIoU}_{\text{Vehicle}}$ $\text{AP}_{\text{Car}}$ $\text{AP}_{\text{Truck}}$
Oracle 72.2 34.6 47.0 21.4
BEVGen [65] 50.1 (-21.1) 5.9 (-28.7) 24.7 (-22.3) 9.1 (-15.0)
MagicD. [18] 58.6 (-13.6) 29.5 (-5.1) 37.3 (-9.7) 17.3 (-4.1)
SimGen-nuSc 60.6 (-11.6) 29.9 (-4.7) 39.1 (-7.9) 18.1 (-3.3)
SimGen 62.9 (-9.3) 31.2 (-3.4) 41.0 (-6.0) 19.6 (-1.8)
Table 5: Comparison involving data augmentation using synthetic data. The Baseline is a single-frame version of BEVFusion [38] trained on nuScenes train set.
Method Map Seg Object Det
$\text{mIoU}_{\text{Road}}$ $\text{mIoU}_{\text{Vehi}}$ $\text{AP}_{\text{Car}}$ $\text{AP}_{\text{Truck}}$
Baseline 72.2 34.6 47.0 21.4
BEVGen [65] 71.9 34.2 47.3 21.1
MagicD. [18] 77.4 37.7 48.0 22.8
SimGen-nuSc 77.7 38.0 48.3 23.0
SimGen 78.9 39.0 49.1 23.6
4.2 Ablation Study
Table 6: Ablation on designs in SimGen. All proposed designs contribute to the final performance.
[…]
Baseline 19.5 45.7
+ Cascade Pipeline 17.2 46.3
+ ExtraCond 17.7 47.6
+ Unified Adapter 16.9 48.2
The ablation is conducted by training each variant of our model on a DIVA subset with 30K frames, and we report FID and average precision of cars ($\text{AP}_{\text{Car}}$) as the quality and controllability metrics. We gradually introduce our proposed components and conditions, starting with a ControlNet baseline [79] that directly takes SimCond as input. As shown in Sec. 4.2, by introducing a cascade pipeline to transform SimCond into RealCond, the FID significantly reduces by -2.3, as the transformed conditions closely resemble real scenarios. Including simulator-pulled ExtraCond to the control conditions improves the alignment of the generated images with the target layouts, effectively enhancing the $\text{AP}_{\text{car}}$ by +1.3. However, a slight deterioration in the FID metric (+0.5) may result from condition conflicts. Lastly, using a Unified Adapter helps alleviate conflicts, significantly improving generated image quality by -0.8. The effectiveness of each modality in ExtraCond is exhibited in Fig. 6, where the addition of instance map, rendered RGB, and top-down view enables the model to better handle object boundaries, orientation angles, and occlusions.
Figure 5: Generating safety-critical scenes. SimGen can also recreate image sequences of safety-critical scenes where risky driving behaviors like sudden braking and merging happen.
Figure 6: Ablation study of simulator conditions.
4.3 Discussions
Extension to video generation. SimGen is not designed for video generation. But the high-quality image generation brings a potential for video generation, which is important for interactive scene generation and closed-loop planning. We have a preliminary attempt by integrating temporal attention layers into UNet similar to [76], and then conducting subsequent training stages focusing solely on learning the newly added layers while freezing the original parameters. This shows a promising result of temporal consistency across frames, as compared with video generation models in Sec. D.2.
Generating safety-critical scenarios. The key innovation of SimGen is the controllability of layouts brought by connecting to a driving simulator. Building upon video generation, we showcase SimGen’s generalization capabilities in novel layouts, specifically in safety-critical scenarios in Fig. 6. The visualized layout is initialized from a scenario sampled from the Waymo Open dataset [63] and then populated with risky behaviors via an adversarial interaction traffic flow generation method [78]. SimGen can transform safety-critical driving scenarios from the simulator into realistic sequential images, including risky behaviors like sudden braking, crossroad meeting, merging, sharp turning, etc. This application is impossible with existing models, which are only trained and conditioned on a given static real-world dataset that lacks records of dangerous driving behaviors. This brings new opportunities for closed-loop data generation capabilities (Sec. D.2).
5 Related Work
Diffusion-based generative Models. Diffusion models have made significant strides in image generation [13, 43, 47, 49, 55, 60] and video generation [4, 22]. Recent works incorporate additional control signals beyond text prompts [21, 37, 45]. ControlNet [79] integrates a trainable copy of the SD encoder for control signals. Studies like Uni-ControlNet [81] and UniControl [52] have also focused on fusing multimodal inputs into a unified control condition using input-level adapter structures. Our method distinguishes itself in its capability of multimodal conditioned generation by addressing the sim-to-real gap and condition conflicts in the complex realm of driving scenarios.
Controllable generation for autonomous driving. Autonomous driving research heavily relies on paired data and layout ground truths, spurring numerous studies on their generation [10, 42]. Some works [19, 26, 76] utilize diffusion models to generate future driving scenes based on historical information, but they lack the ability to control scenes through layout. Other generative methods, like BEVGen [65] and BEVControl [77], use BEV layouts to create synthetic single or multi-view images. Recent innovative method Panacea [70] generates panoramic and controllable videos, while MagicDrive [18] offers diverse 3D controls and tailored encoding strategies. Lastly, DriveDreamer [68] and DrivingDiffusion [36] employ diffusion models for realistic multi-view video generation and environment representation. Yet, these works are confined to limited appearances and layouts of static datasets, restraining their real-world applicability and the controllability over the layouts that deviate from the dataset, such as the safety-critical scenarios.
Scenario generation via simulators. Driving simulators [14, 34] are fundamental to autonomous driving development, providing controlled simulations that mimic reality. Notable studies include SYNTHIA [59], AIODrive [71], and GTA-V [56] that generate vitural images and annotations. SHIFT [64] diversifies with environmental changes, while CAT [78] creates safety-critical scenarios for targeted training from real-world logs. Despite their layout diversity and attempts at photorealism enhancement [57], the simulated images lack realism. In this work, we bridge the two worlds to obtain both the appearance diversity from diffusion models and the layout controllability from simulators.
6 Conclusion
We propose a simulator-conditioned diffusion model, SimGen, that learns to generate diverse driving scenarios by mixing data from the simulator and the real world. A novel dataset containing massive web and simulated driving videos is collected to ensure diverse scene generation and mitigate simulation-to-reality gap. By obtaining diversity in appearance and layout, SimGen exhibits superior data augmentation and zero-shot generalization capabilities in generating diverse and novel scenes.
Limitations and future work. SimGen currently does not support multi-view generation, limiting its application in Bird’s Eye View models. Inheriting the drawbacks of diffusion models, SimGen suffers from long inference time, which may impact the applications like closed-loop training. The study of extending SimGen to video generation is left for future work.
Appendix
[…]
1
Appendix A Discussions
SimGen project page provide links to the YouTube videos (DIVA-Real_Video_Links) used in DIVA-Real, as well as digital twins of nuScenes dataset (DIVA-Sim_nuSc_Digital_Twins) and safety-critical video clips (DIVA-Sim_Safety-critical_Demo_Videos) included in DIVA-Sim.
To better understand our work, we supplement with the following question-answering.
Q1. What makes SimGen stand out compared to pixel-to-pixel transformation models?
Recent GAN-based and Diffusion-based works in image transformation can generate images that are controllable based on specific conditions [21, 60]. Yet, their limitations lie in the fact that the content they generate is strictly tethered to these input conditions. If these conditions, derived directly from a simulator, are missing important contextual information like backgrounds and buildings, then output images may similarly lack background details Consequently, SimGen employs a cascade structure, permitting the CondDiff model to conceptualize different background scenarios through text, thereby enriching the visual composition of the rendered driving scenes. Detailed analysis is shown in Sec. C.1.
Q2. What is the criteria to demonstrate good generalization and diversity of your model? How much data do we need?
Currently, it’s challenging to define a specific standard to assess the diversity and generalization abilities of the models, as quality evaluation is subjective and fair comparison can be difficult. However, by utilizing publicly available data, we have found that scaling up the data size proves beneficial for the zero-shot generation on novel scenarios. Equally important to note is that our approach is easily scalable, and by leveraging massive in-the-wild data, we offer a continuing opportunity to strengthen its generalization capabilities.
Q3. Broader impact. What are potential applications and future directions with the provided DIVA data and the SimGen model, for both academia and industry?
Datasets. DIVA collects massive data from YouTube and simulators, significantly enhancing the appearance and layout diversity of driving video clips. This provides the community with extensive high-quality resources for exploring open avenues in autonomous driving and Sim2Real research.
Models. Beyond data augmentation, we hope our model can also benefit the community by enabling wider applications. In this work, we demonstrate SimGen’s capability as a closed-loop data generator. It holds promise to adapt to downstream tasks like closed-loop evaluation of autonomous driving agents [48], which is showcased in Sec. D.2. To boost deployment efficiency, distilling knowledge from the generative model is worth exploring [32]. Besides, simulator-conditioned scene generation also provides opportunities to achieve physically grounded real-world generation [2, 15]. Please note that our model will be publicly released to benefit the community and can be further fine-tuned flexibly according to custom data within the industry.
Negative societal impacts. The potential downside of SimGen could be its unintended use in generating counterfeit driving scenarios. Our code includes the Diffusers [67] safety checker to screen for NSFW outputs. Besides, we plan to regulate the effective use of the model and mitigate possible societal impacts through gated model releases and monitoring mechanisms for misuse.
Q4. Limitations. What are the issues with current designs, and corresponding preliminary solutions?
1) Panoramic image generation is necessary for current Bird’s Eye View perception models in autonomous driving. Yet, to utilize scaled web data, which consists of front-facing single-camera footage, SimGen does not engage in multi-view image generation. This may limit the application of SimGen in real-world deployments. 2) We chose SD-2.1 [58] as our base diffusion model, inheriting its advantages of high visual quality and better rendering capabilities of the text encoder. On the other hand, we noted that it has a slow sampling speed and high computation costs. Our model does indeed suffer from this issue.
However, as a pioneering work exploring how to introduce simulators into generative models for diverse driving scene generation, the primary focus of this work is the simulator-conditioned generalization ability across unseen driving scenarios rather than multi-view designs and computational overhead. Future work might include trying cross-frame attentions [18, 36, 70], faster sampling methods [44, 61, 82], and transferring our general method to more efficient diffusion models.
Appendix B DIVA Dataset
Our dataset, DIVA, contains 147.5 hours of driving video along with diverse multimodal conditions, including text, segmentation, depth, and virtual images. In this section, we detail the YouTube and simulator video collection process, annotation method, more examples, and analysis to illustrate the diversity of the DIVA dataset.
B.1 DIVA-Real
B.1.1 Data Collection and Cleaning
Data preparation. We first searched for driving videos on YouTube using keywords such as driving videos, 4K, and HD. We then identified a selection of YouTubers who consistently upload high-quality driving videos. We further inspected the quality of these videos in terms of resolution, resulting in 130 high-quality front-view driving videos, including Barcelona 4K - Driving Downtown, Cairo 4K - Pyramid Expressway Sunrise, Las Vegas 4K - Sunset Drive and Istanbul 4K - Night Drive - Turkey. We used videos from 10 selected clips as the validation set and the other videos for training. The diversity of DIVA-Real is illustrated in Fig. 7. Additionally, we cut off the first and last 30 seconds of each video to remove any solicitations or other edited footage.
Data format. We segmented the video data into images at a rate of 10Hz, with a resolution of 1080p (1960$\times$1080) for each image.
Data cleaning. To automate the process of filtering out low-quality images from the dataset, we utilize a vision language model (VLM), LLaMA-Adapter V2 [17]. First, we group the images at a rate of 2Hz and randomly select one image from each group to feed into the VLM. We provide a checklist, asking the VLM sequenced questions about the image quality. The checklist includes items such as non-front view, video transition, black screens, etc. The VLM then uses its acquired world knowledge to infer and assist in automatically eliminating low-quality images. The checklist is organized as a set of text prompts given to the VLM, specified below.
Text Prompt Examples: "Is the driving scenario image presented from a POV or other perspective rather than a front view?", "Does the driving scenario image exhibit a gradual transition due to a video transition?", "Is the image almost completely black, distinguishable from those depicting night driving?", "Is this image excessively blurry, rendering any foreground object information indistinguishable?", "Does this image feature subtitles, distinct from signs or markers in driving scenarios?", "Does this image depict a scenic view or a bird’s-eye perspective, distinct from front-view driving footage?", etc.
Figure 7: Various video samples from DIVA-Real. Due to space limitations, we only showcase certain frames from the videos. It covers a wide range of diversity across multiple axes, including geographical location, traffic scenarios, time periods, weather conditions, etc.
B.1.2 Multimodal Annotation
Our OpenDV-Wild features three types of annotations: text, depth, and semantics. We leverage the established BLIP2-flant5 [33] to describe each frame’s main objects or scenarios with the following prompt.
Text Prompt Example: "Question: Describe the image of a driving scenario concisely. Answer:".
We present a content query to the VLM according to each example’s text prompts. If the VLM responds with no to all the queries, then the set of images can successfully pass the review.
For depth and semantic segmentation, we employ pre-trained ZoEDepth [3] and Segformer [74] for automated label generation. Segformer is previously trained on the CityScapes dataset [12]. The examples of text annotations are shown as follows.
Text Annotation Examples: "A car is driving down a street in Rio De Janeiro.", "A motorcyclist on a road with a speed limit sign.", "A car driving down a street in Madrid, Spain.", "A view of a city street with tall buildings and a TV tower in the background.", "A snowy road with trees on both sides of the road and a red car is driving.", "A car driving on a highway at dusk in Las Vegas, Nevada.", etc.
Table 7: Location distribution of DIVA-Real.
Continent North America South America Europe Asian Africa
Proportion 56.9% 8.5% 16.9% 14.6% 3.1%
Table 8: Time period distribution of DIVA-Real.
Period Daytime Dawn Dusk Nighttime
Proportion 55.8% 16.3% 10.1% 17.8%
Table 9: Weather distribution of DIVA-Real.
Weather Normal Rainy Cloudy Foggy Snowy
Proportion 58.2% 1.0% 28.6% 2.1% 10.2%
B.1.3 Appearance Diversity Highlights
Diversity over prior datasets. Beyond its large data scale, our dataset outshines competitors in terms of appearance diversity. YouTube, known for its diverse content, is regarded as a global mosaic and a crucial source of data. DIVA-Real leverages this by comprising 120 hours of publicly available videos from over 71 cities across more than 19 countries. Our dataset exhibits a globe-wise geographic distribution compared to other open datasets collected in limited regions. On top of that, DIVA-Real covers a rich variety of driving scenarios, including bridges, bays, wilderness, deserts, dusk, fog, and more. Unlike datasets restricted to dull scenes, our dataset empowers models to capture a wealth of visual appearance diversity. Lastly, ours boasts a richness of annotations on par with other datasets.
In this section, we provide a detailed data analysis about the diversity of DIVA-Real. For simplicity’s sake, we assume that all clips within a video are shot at the same location and time, with any single frame from a segment representative of the geographical location, lighting conditions, weather, and other informational aspects concerning the video. Thus, we manually review each video’s title and a random frame from the video and assess its geographic location, time, and weather conditions, among other things, for statistical analysis. The following diversity analysis has been derived from this process in three aspects.
Location distribution. According to statistical results, YouTube videos are derived from 71 cities in 19 countries, covering a much larger area than any existing public driving dataset, as shown in Tab. 7. For example, in the most popular regions, DIVA-Real includes 67 hours of video data in the United States, covering cities like Los Angeles, New York, Las Vegas, Miami, Boston, Atlanta, New Orleans, etc., and encompasses geographical areas such as urban, rural, coastal, wilderness, mountainous, and port regions.
Time period and weather variance. The DIVA-Real dataset also includes a variety of times and weather conditions. As shown in Tab. 8, in addition to daytime, the dataset covers a considerable proportion of dawn, dusk, and nighttime scenarios. Tab. 9 presents the weather distribution in the dataset, including rainy, cloudy, foggy, and snowy conditions. These diverse times and weather conditions ensure a variety of appearances.
Corner cases. YouTube videos also include extreme cases and safety-critical scenarios. Fig. 7 presents several special cases from DIVA-Real, such as crowded pedestrian-filled intersections (fourth one in the first row), roads with a lot of parked vehicles (fifth one in the first row), a country road in the sunset (third one in the fifth row), and passing under an overhead structure with limited light (fifth one in the seventh row).
B.2 DIVA-Sim
B.2.1 Data Collection
The DIVA-Sim dataset is collected through the MetaDrive simulator [34]. DIVA-Sim accumulated a total of 27.5 hours of virtual driving data, including 5.5 hours from the digital twins of the nuScenes dataset [6] via ScenarioNet [35], and 22 hours of dangerous driving scenarios collected initially based on the Waymo Open dataset [63] by adversarial interventions [78]. For data from nuScenes and Waymo Open, each video lasts 20 seconds and 8 seconds, respectively.
For each scenario, DIVA-Sim provides a variety of labels, including rendered RGB, depth, segmentation, instance map, and top-down view. All camera outputs are synthesized using the OpenGL rendering backend from the Panda3D game engine, allowing us to incorporate depth map and semantic colormap similar to ZoeDepth [3] and Segformer [74].
Data format. We segmented the video data into images at a rate of 10Hz, with a resolution of 1960$\times$1080 for each image.
B.2.2 Layout Diversity Highlights
We randomly sample 500 safety-critical videos generated based on Waymo Open dataset [63] and manually reviewed the contents of the top-down view of each video, collecting the data shown in Tab. 10. Beyond forwarding, turning, and stopping, DIVA-Sim also covers cases like changing lanes, passing through intersections, and making U-turns. Fig. 8 visually displays the top-down views of various dangerous driving scenarios, including collisions, quick stops, and reckless merging. These illustrate the diversity of layouts in DIVA-Sim.
Table 10: Layout distribution of DIVA-Sim.
Behavior Forward Left Right Left Lane Right Lane Intersection U-Turn Stop
Turn Turn Change Change Passing
Proportion 36.2% 14.3% 10.0% 10.7% 17.4% 6.6% 1.2% 3.6%
B.3 Extensions on Public Dataset
In addition to collecting YouTube data and simulation data, we also annotate data of public datasets to promote research from simulation to reality and to enable fair comparisons.
nuScenes dataset. The nuScenes dataset [6] is a public driving dataset that includes 1000 scenes from Boston and Singapore for diverse driving tasks [85, 40, 39]. Each scene comprises a 20-second video, approximately 40 frames. It provides 700 training scenes, 150 validation scenes, and 150 test scenes. Similarly, we utilize BLIP2-flant5, ZoeDepth, and Segformer to provide textual, depth, and semantic labels for this dataset.
Figure 8: Various safety-critical layouts from DIVA-Sim. The yellow rectangular represents the ego-car and the other cars interacting with it. All scenarios are initialized by the Waymo Open dataset [63] and generated by adversarial interactions [78] within simulators.
B.4 License and Privacy Considerations
All the data is under the CC BY-NC-SA 4.0 license [^1] https://creativecommons.org/licenses/by-nc-sa/4.0/deed.en. Other datasets (including nuScenes [51], Waymo Open [63], Metadrive [34]) inherit their own distribution licenses.
We place a high value on license and privacy protection, following the precedent from YouTube-8M [1], YouTube-VOS [75], AOC [80], CelebV-HQ [89], ELM [88], OpenScene [11], and Kinetics [28], etc. For videos from YouTube, permission to access the video content is received through a Creative Commons license. Besides, we skip channel-related content at the beginning and end of the videos during data processing to ensure we do not infringe upon the rights of logos, channel owner information, or other copyrighted materials. We do not provide video content; users are redirected to original YouTube videos via a link. The platform safeguards personal info with encryption, access limits, and identity checks to prevent unauthorized video access. We will credit the source, provide a link to the license, and state that no modifications have been made to the video itself, and the data will not be used for commercial purposes. All the data we obtain complies with regulations and YouTube’s Privacy Policy. In addition, we comply with any limitations required by applicable law and any requests submitted by users. For instance, users may have the right to view, correct, and delete personal information we possess about them, such as deleting text labels, unlinking videos, and de-identifying data.
Appendix C Implementation Details of SimGen
C.1 Empirical Study
C.1.1 Cascade Diffusion Scheme
Figure 9: Gaps between conditions in simulators and reality. In each group of images, the first row represents the conditions in the simulator, and the second row represents the real conditions.
It’s worth noting that, despite our efforts to replicate real scenes as closely as possible in the simulator, gaps are inevitably present. Gaps between the conditions in simulators and reality can affect the accuracy of model training. These gaps primarily originate from: 1) mismatches, 2) inherent flaws of 3D models, and 3) missing backgrounds. In Fig. 9, we provide some visualizations to illustrate these gaps.
The mismatch between simulated and real images can exist in multiple aspects. Firstly, given that the models within the simulator are finite, it is unrealistic to accurately represent a wide variety of object categories through it. As shown in example 1.a, for excavators, the simulator uses dump trucks based on size, resulting in some mismatches in object shape. Secondly, slight differences between the camera positions in the simulator and real datasets can cause objects to be displaced in the simulator, as in example 1.b, where $\Delta y_{1}$ is larger than $\Delta y_{2}$. Thirdly, the simulator lacks physical entities such as buildings or trees, making it difficult to replicate scenarios where vehicles are obstructed by the environment, as shown in example 1.c. Beyond mismatches, inherent flaws may also exist within the simulator models. Although the simulator can reflect the real size of vehicles, achieving perfect consistency in shape is still challenging. Furthermore, the transparency of the model’s window parts can lead to discrepancies between depth and semantic segmentation results in those areas. Finally, the simulator lacks elements such as buildings, trees, and street lamps, which are essential conditions for controlling image generation.
Through the above analysis and visualization results, combined with the outcomes of our ablation study, we infer that real depth and semantic segmentation images are more suited to control real image generation than simulated equivalents. Therefore, SimGen’s cascade diffusion network is necessary, as it allows the decoupling of the introduction of SimCond in the CondDiff module from the controlled image generation in the ImgDiff module.
It’s worth noting that this gap doesn’t mean simulated conditions can’t be used for training image generation models. Instead, it suggests that employing real depth and semantics is a better choice, and better aligned for joint training with data from YouTube. Indeed, the simulator’s ExtraCond can also contribute to the model’s accuracy. For instance, the instance map can indicate the object count and occlusion relationships in the scene to the model, and the top-down view can provide spatial location information. These insights can guide the use of a unified adapter to merge multi-modal conditions.
C.1.2 Unified Adapter
SimGen employs a unified adapter to address the two obstacles in multi-modal condition conflicts: Modal Discrepancy and Condition Disparity. Modal discrepancy refers to the inconsistent number of modalities between data from nuScenes and YouTube (the latter lacks ExtraCond), which might lead the model to establish a statistical shortcut, such as outputting nuScenes-style images when ExtraCond is present and YouTube-style images when it’s absent. This shortcut can significantly impact the model’s instruct-following ability and diversity at inference time. On the other hand, condition disparity refers to the lack of background information in the ExtraCond condition, which can result in conflicting control information with RealCond. Thus, our proposed solution is to use an adapter to merge various modalities into a unified control feature, employing a mask during the fusion process to eliminate conflicts arising from absent background information in ExtraCond.
C.2 Model Design
Realism-controllability trade-off. Apart from the discretization steps of the CondDiff solver, the critical hyperparameter for sim-to-real transformation is $t_{s}$, the starting time of the image synthesis process in the reverse SDE. We notice that with a fixed CondDiff model, there’s a trade-off between Realism and Controllability when choosing different $t_{s}$ values. Smaller $t_{s}$ values lead to fewer denoising steps, giving SimCond more control over image generation but potentially compromising realism. Generally, we find $t_{s}\in[0.4,0.65]$ to work well for the foreground, and we ultimately select $t_{s}$ as 0.5 for foregrounds. Then, we mask the foreground part, and there will be an additional $t_{s}=0.5$ for background generation.
Extension on Video Generation Having acquired the single-frame variant, we lock the original blocks within the denoising UNet and intersperse them with temporal reasoning blocks, mirroring the strategy of GenAD [76], thereby facilitating video sequence modeling.
C.3 Training Details
SimGen is trained in two modules: CondDiff, which converts simulated conditions to real ones, and ImgDiff, which generates images from multimodal conditions. In the first stage, we fine-tune the pre-trained SD-2.1-V on per-image denoising with 1.1B trainable parameters of its denoising UNet. It is trained on 4.5M text-depth-segmentation pairs of DIVA-Real and nuScenes. We train the model for 30K iterations on 8 GPUs with a batch size of 96 with AdamW [41]. We linearly warm up the learning rate for $10^{3}$ steps in the beginning, then keep it constant at $1\times 10^{-5}$. The default GPUs in most of our experiments are NVIDIA Tesla A6000 devices unless otherwise specified.
In the second stage, we train the model via a unified adapter and ControlNet using text-condition-image pairs, lifting it to generate realistic images during inference. The training data consists of DIVA and nuScenes, with conditions confined to ExtraCond and RealCond. Following the design of ControlNet, we freeze the input and middle layers of the UNet, training only the parameters of the control branch and the Adapter. This stage is trained for 50,000 iterations on 8 GPUs, with a 295 batch size of 96.
For the extension of video generation, we freeze all blocks of the single-frame version and only optimize the introduced temporal reasoning blocks, resulting in 418M trainable parameters in this stage. To maximize the data efficiency for constructing video clips, we take each frame of a 10Hz YouTube and nuScenes video as a starting frame to form a 3s training sequence at 2Hz. The text condition is structured in the same way as the first stage, and we acquire the context from the middle frame of the sequence. SimGen is trained on 8 GPUs for 30K iterations with a total batch size of 24. The learning rate is set as $1\times 10^{-5}$ after $10^{3}$ warm-up steps.
In both stages, the input frames are resized to $256\times 448$, and the text condition $c$ is dropped at a probability of $\gamma_{c}=0.1$ to enable classifier-free guidance [24] in sampling. Both CLIP text encoders and the autoencoder are kept frozen throughout our experiments.
For effective classifier-free guidance [24], ImgDiff random drops conditions during training at a rate of $\gamma_{c}=0.1$. Additionally, we enhance the model’s robustness by randomly masking the background of real conditions with a fixed probability of $\gamma_{b}=0.5$. To address potential cumulative errors from the CondDiff process during ImgDiff denosing, we introduce slice noise with a probability of $\gamma_{n}=0.25$. Slice noise entails partitioning the image into $n\times n$ patches and randomly masking them with a probability of $\gamma_{p}=0.25$.
C.4 Sampling Details
Given conditions from simulators, SimGen first has a reverse SDE process in CondDiff. It starts from SimCond added with standard Gaussian noises. The sampling step of this stage is 25 ($t_{s}=0.5$). After that, the second sampling process is involved in ImgDiff, starting with random Gaussian noises. Both sampling processes are performed by Denoising Diffusion Implicit Models (DDIM) [62]. We use 50 sampling steps and set the scale of classifier-free guidance to 9.5. The sampling speed is 1.13 seconds per step per batch. The image resolution is $256\times 448$, and the video sequence is at 2Hz.
Appendix D Experiments
We conduct extensive experiments on multiple datasets to evaluate the performance of our method. For comparison convenience, we trained two models on the nuScenes and DIVA datasets, respectively, namely SimGen-nuSc and SimGen, adopting the same training strategy.
D.1 Metrics
One of the roles of synthesized images is to augment existing perception models of autonomous driving [83, 84, 87, 86]. We use various metrics in multiple aspects for quantitative evaluation. For generation quality metrics, we use Fréchet Inception Distance (FID) [23] and Fréchet Video Distance (FVD) [66]. For generation diversity, the pixel diversity $D_{\text{pix}}$ metric is leveraged. The controllability of the model is reflected by the alignment between the generated image and the conditioned BEV sequences. For the video generation task, all frames are at 2Hz. The specific metrics are described as follows.
FID: It evaluates the generation quality of images, which are video frames in our experiments, by measuring the distribution distance of features between the predictions and original frames in the dataset. The features are extracted by a pre-trained Inception model. For quantitative comparison on nuScenes, FID is evaluated on 6019 generated frames and ground-truth frames. For experiments on YouTube, FID is calculated on 18000 frames from both generation and the dataset.
FVD: It measures the semantic similarity between real and synthesized videos with a pre-trained I3D action classification model [8] as the feature extractor. We evaluate 4369 video clips from nuScenes and 3000 video clips from YouTube.
$D_{\text{pix}}$: To gauge the diversity of the generated data, we compute the standard deviation of the pixel values in the generated images. A higher value indicates a greater diversity of colors in the generated data. The conditions for the evaluation come from the nuScenes validation set, and text prompts are collected from both the nuScnes validation set and randomly selected 18,000 frames of YouTube data. For each condition, we randomly select a text from the collected prompts as input, and the model generates the corresponding image. To reduce randomness, we test each model 3 times with the same random seed and take the average.
Figure 10: Comparison of proposed cascade diffusion model (blue boxes) to naïve approaches (gray boxes).
Figure 11: Visualization of text-grounded Sim-to-Real condition transformation.
Figure 12: Preliminary attempt at closed-loop evaluation. For two scenarios, the IDM behavior [34] (gray boxes) leads to hazardous driving situations, while manual control (blue boxes) adopts measures to evade risks.
Controllability Metrics: To affirm the consistency between the generated images and original data in layout, we employ metrics such as Average Precision (AP), and Mean Intersection over Union (mIoU) to evaluate the perception performance on the nuScenes dataset. Our evaluation comprises two aspects: firstly, we use pre-trained perception models to compare the validating performance of generated data with real data. Secondly, we explore the potential of employing augmented training sets as a strategy for performance improvement.
We adopt BEVFusion [38], a state-of-the-art perception method, as our primary evaluation tool. Specifically, we utilize a BEVFusion implementation that only incorporates a front view, masking other camera perspectives and ground truth during evaluation.
D.2 More Ablations
Table 11: Quality of video generation.
[…]
DriveGAN [30] 502
DriveDreamer [68] 452
DrivingDiffusion [36] 332
GenAD [76] 184
SimGen 271
Results of video generation. It’s worth noting that the innovation of SimGen doesn’t focus on video generation. However, high-quality image generation implies the potential for video generation, which is crucial for interactive scenario generation and closed-loop planning. We made a preliminary attempt at video generation based on GenAD [76]. Sec. D.2 compares SimGen with other video generation models. Thanks to its commendable image generation quality, SimGen achieves performance that is on par with other models.
Effectiveness of Sim2Real condition transformation. To validate the effectiveness of the cascade diffusion model, we display a set of comparative images in Fig. 10. The blue boxes represent the cascade diffusion structure we adopt, it transforms SimCond from the simulator into RealCond in CondDiff, and then generates realistic images through the ImgDiff model. The grey boxes signify that we removed CondDiff, and ImgDiff directly generates images using SimCond. It is observable that removing CondDiff causes the generative model to introduce certain distortions from the simulator conditions into the produced images, as seen in the overly narrow wheels and deformed rear part shown in the left set of pictures. The introduction of CondDiff transforms these distortions into realistic conditions after resampling, thus greatly enhancing image generation quality. Fig. 11 further demonstrates that CondDiff can transform depth and segmentation from the simulator into real ones based on different texts.
Closed-loop evaluation. We further explore applying our simulator-conditioned generative models to the closed-loop evaluation in Fig. 12. The evaluation focuses on two driving behaviors, namely IDM [34] (gray boxes) and manual control (blue boxes) in different scenarios. IDM could lead to risks like sudden braking or collision in these cases. Conversely, manual control promotes safety by maintaining distance and slowing down. The video data is generated by SimGen, using conditions pulled from simulator interactions, with a one-second frame interval.
Generalization on novel simulators. As CondDiff can convert simulated conditions into real conditions in an adaptation-free approach, SimGen possesses the ability to perform zero-shot generalization to other simulators. In Fig. 13, we exhibit a case study of generating realistic images using depth and semantic segmentation conditions provided by CARLA [14]. This provides the possibility for SimGen to utilize and integrate the diverse layouts, driving policies, and physical engines provided by various simulation platforms to generate diverse driving scenarios.
D.3 Qualitative Results
Text-grounded image generation. SimGen is a capable text-to-image diffusion model for driving scenarios, especially when examining text controllability compared to other works. In Fig. 14, we demonstrate SimGen’s exceptional ability to generate images from different text prompts. Thanks to the relatively simple simulator conditions and comprehensive DIVA dataset, the text prompt can effectively influence the resulting image, even changing the surrounding building and background with reference to specific cities. Whereas as other diffusion-based generative models struggle to generate images with characteristics that were not originally present in nuScenes  [18], like unseen weather or background settings, Simgen’s text-grounding can influence both foreground objects like cars and match the background to cities not present in nuScenes.
Simulator-conditioned image generation. Fig. 15 displays additional examples of SimGen generating diverse images based on conditions provided by the simulator. The far-left column presents the simulator-rendered RGB, while the right side shows images generated by SimGen following different text prompts. This further validates SimGen’s potent ability to adhere to the simulator conditions while maintaining rich appearance diversity.
Video generation. Our preliminary SimGen video generation model is able to maintain the ability to create driving images with a wide range of backgrounds while also incorporating temporal consistency as shown in Fig. 16.
D.4 Failure Cases
We show some failure cases of SimGen in Fig. 17. The first column represents the conditions from the simulator, with each generated image accompanied by a text prompt. The failure cases of SimGen are included as follows: 1) Text comprehension error: as in the first image, where the adjective "green" is not assigned to any discernible "traffic light" but instead as a vehicle color. 2) Condition conflict: the second image provides a text prompt of a tow truck, but it’s challenging for the model to generate such a vehicle based on the shape of a sedan. 3) Background subsumption: the third image demonstrates a case where the background subsumes the smaller car. 4) Generation instability: in the last image, SimGen occasionally produces distortion and blur in background generation, likely due to cumulative model error and overabundance of nighttime images in DIVA.
Figure 13: Case study of zero-shot image generation on CARLA. We randomly select a scenario in CARLA, for which SimGen generates various driving scenarios through the conditions (depth and segmentation) produced by the simulator and different textual prompts.
Figure 14: Text-grounded image generation. Each image is generated by SimGen using a randomly selected text prompt and simulator conditions. The rich appearance diversity is reflected through the wide range of generated content.
Figure 15: Diverse generated images from simulator scenarios From the same original simulated driving scenario (left column), we show a diverse range of generated images (columns 2 through 7). SimGen is capable of generating driving scenes in a wide variety of settings based on the same simulator conditions.
Figure 16: Preliminary attempt at video generation. Notably, SimGen is not designed for video generation. We simply follow some practices in [76] to temporal consistency, and video generation will be our future work.
Figure 17: Failure cases of SimGen.